OpenStudy (anonymous):

(sinx-cosx)squared=1+2 sinxcosx

OpenStudy (anonymous):

1-2sinxcosx = 1+2sinxcosx -sinxcosx = sinxcosx sinxcosx = 0 \[\sin(x) = 0 or \cos(x) = 0\] x=1/2n*pi, \[n \in \mathbb{R} \]

OpenStudy (anonymous):

how did u do it from the beginning? should this become an identity or not?

OpenStudy (anonymous):

? this is not an identity. To simplify this let sin x = a; cos x = b; a^2+b^2 = 1; (a-b)^2 = a^2+b^2 - 2ab

OpenStudy (anonymous):

so for sinx and cosx, when(sinx-cosx)^2 is used, how did u use 1 and 2?

OpenStudy (anonymous):

let sin x = a; cos x = b; (a-b)^2 = (a-b)*(a-b) = a*a - a*b - b*a + b*b = a^2+b^2 - 2ab a^2+b^2 = 1 (sin^2 x + cos^2 x = 1) = 1 - 2ab

OpenStudy (anonymous):

oh i get it! thanks! can u help me with the other question i posted up a few minutes ago?

OpenStudy (anonymous):

its cosx/secxsinx=cscx-sinx

OpenStudy (solomonzelman):

This is trig and you need to find x, let me try to help you on this one!

OpenStudy (solomonzelman):